If 10 kg/min of air are compressed isothermally from P1=96 kPa and V1 =7.65 m^ 3/min to P2 =620 kPa , find the work, the change of entropy and the heat for (a) non flow process and b) a steady flow process with v1= 15 m/s and v2 =60 m/s .
1
Expert's answer
2021-01-11T07:41:44-0500

Solution

a) for non flow process

This is isothermal process so

Change in internal energy

"\Delta U=0"

"P_1V_1=nT_1\\T_1=\frac{P_1V_1}{n}"

Where n=0.3452Kg

After putting value of pressures and n then

"T_1=255.9K"


n = m / M

= ( 10 kg )/ ( 28.97 kg/kgmol )

= 0.3452 kgmol


Work done

W=nRT"\ln{\frac{P_1}{P_2}}"

W=0.3452"\times" 8.314"\times" 255.9"\ln{\frac{96KPa}{620KPa}}
\\"

W=-1370KJ

Now heat

"Q=\Delta U+\Delta W\\Q=-1370KJ"

Now entropy change

"\Delta S=\frac{Q}{T}=\frac{-1370KJ}{255.9}"

"\Delta S=-5.354KJ/kelvin"

b) now for steady flow

Work done

Ws=0.3452"\times" 8.314"\times" 255.9"\ln{\frac{96KPa}{620KPa}}
\\"

"W_s=-1370KJ\/min"

"\Delta H=0"

"\Delta KE=\frac{10\times(60^2-15^2) }{2}=14" KJ/min

"\Delta PE=0"

Total heat

"Q=\Delta H+\Delta KE+\Delta PE+" Ws

Q=-1353KJ/min

Now entropy change

"\Delta S=\frac{Q}{T}=\frac{-1353KJ/min}{255.9Kelvin}"

"\Delta S=-5.267" KJ/min-Kelvin